OpenStudy (anonymous):

I have a logarithmic problem solve for x approximate 4^(x+4)=5^(x-6)

OpenStudy (anonymous):

take logs both sides ( doesnt matter what base, as long as they are the same )

OpenStudy (anonymous):

ill use base e

OpenStudy (anonymous):

take log on both side, then you'll get (x+4)log 4 = (x-6)log 5, solve it and get the answer.

OpenStudy (anonymous):

ln [ 4^(x+4) ] = ln [ 5^(x-6) ] now use log laws to bring power down in front

OpenStudy (anonymous):

etc ...

OpenStudy (anonymous):

I don't come up with the correct answer I have had it wrong 5 times

OpenStudy (anonymous):

I have 4 log 5+4log5 / log5-log 4 how do I solve it with the calculator they come up with an approximate

OpenStudy (anonymous):

I don't have my calculator right now with me, nor my internet connection is working fine and this site is damn slow or else i would have solved that for you.

OpenStudy (anonymous):

how do you solve it

OpenStudy (anonymous):

it's roughly 68.1257..

OpenStudy (anonymous):

how did you get that

OpenStudy (anonymous):

I have another one you might be able to help me with use log b 2=0.693 and /or lo b 7 =1.946 to find log b 14

OpenStudy (anonymous):

the b is below log

OpenStudy (anonymous):

you have to get an approximate also

OpenStudy (anonymous):

don't have any clue how to do it no examples

OpenStudy (anonymous):

I only come on when I am stuck on problems

OpenStudy (anonymous):

\[4^{x+4} = 5^{x-6}\] \[(x+4)\log4 = (x-6)\log6 \] \[x \log(4)+4 \log(4) = x \log(5) - 6 \log (5)\] \[x \log (4) - x \log (5) = -6 \log (5) - 4 \log (4)\] \[x(\log(4)-\log(5)) = -6 \log (5) - 4 \log (4)\] \[x = {{-6 \log (5) - 4 \log (4)} \over {\log (4 ) - \log (5)}} \approx 68.1257\]

OpenStudy (anonymous):

there's a typo in the second line - it's supposed to be: (x+4) log 4 = (x-6) log 5 sry, about that...

OpenStudy (anonymous):

if you'll give me medal then here's the solution: 4^(x+4)=5^(x-6) => (x+4)log 4 = (x-6)log 5 =>(x+4)/(x-6) = log5/log 4 => (x+4)/(x-6) = 1.16 => (x+4) = 1.16 (x-6) =>1.16x -x = 4+1.16 (6) => x = 10.96/0.16 => x = 68.5

OpenStudy (anonymous):

if you solve this one I will give you a medal

OpenStudy (anonymous):

use log b 2=0.693 and /or log b7=1.946 to find log b 14 log b 14 = approximate the b is below log

OpenStudy (anonymous):

\[\log_{b} (2) = 0.693\] \[\log_{b} (7) = 1.946\] like this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay - you just need to remember one simple fact: \[\log_{b}(x) = {\log_{k}(x) \over \log_{k}(b) }

OpenStudy (anonymous):

I have got this one wrong 5 times also

OpenStudy (anonymous):

\[\log_{b}(x) = {\log_{k}(x) \over \log_{k}(b) } \]

OpenStudy (anonymous):

okay what do I plug in where

OpenStudy (anonymous):

I got 2.808

OpenStudy (anonymous):

b is your unknown base and k is just some random base you can choose... for example \[\log_{15} (80) = { \log_{10}(80) \over \log_{10}(15) }\] so you got the base-10 logarithm on your calculator and can therefore calculate teh base-15 logarithm of 80...

OpenStudy (anonymous):

I just don't get it can you help me this please solve I just want to get this BS done it is my last problem

OpenStudy (anonymous):

I have tried so many different ways to solve it and emailed my professor and he is no help

OpenStudy (anonymous):

lol, one should do his/her assignment himself. just kidding, don't mind!!

OpenStudy (anonymous):

okay, okay - BUT, you rly should try to understand it, cause it's quite important. \[\log_{b}(2) = {\ln (2) \over \ln(b)} = 0.693\] \[\ln(b) = {\ln(2) \over 0.693}\] \[b = e ^ {\ln(2) \over 0.693} \approx e ^ 1 \approx 2.7183\] the last line is due to the fact, that ln(2) is about 0.6931... now you should try the same with the other fact you got and check the result.

OpenStudy (anonymous):

okay thank you